Direct Labor Variances Bellingham Company produces a product that requires 4 standard direct labor hours per...

  1. Direct Labor Variances

    Bellingham Company produces a product that requires 4 standard direct labor hours per unit at a standard hourly rate of $20 per hour. If 15,000 units used 61,800 hours at an hourly rate of $19.85 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    a. Direct labor rate variance $
    b. Direct labor time variance $
    c. Direct labor cost variance $
Answer #1

Labor Rate variance = (SR-AR) * AH

= (20-19.85)*61,800

= -9270 Favourable

Labor time variance = (SH-AH) *SR

= (15000*4 - 61800)*20

= 36,000 Unfavorable

Labor cost variance = 9270F + 36000 U

= 26,730 Unfavorable